Shift report 20011109 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Nov 09 2001 - 08:07:31 EST

  • Next message: Apache: "Shift report 20011109 08:00-16:00"

    ------------------------------------------------------------
    Date: Fri Nov  9 08:07:31 2001, Shift: 00:00-08:00
    Supervisor: TML, with: 
    Log book#10 pages 117 - 118
    
    Shift Summary:
    ------------------------------------------------------------
    1. We had 1 long store. Ended with low rate, 1.5 hours 
    after Brahms and Phoenix voted to dump. New Run started 
    just started. 
    
    2. Run   trig1  trig6
       5734  320k   101k  stopped to change D5 to 4B
       5735  162k   50k   stopped to change FS polarity to B
       5736  117k   33k
       5737  just started.
    
    3. Had DC off in 5735 & 5736 hadde map-error flashing 
    every 1-2 min.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Nov 09 2001 - 08:07:56 EST